### v1

@@ 3.1 General Procedure for Using Loss Functions
3.2 Preexisting Outlier Criteria
Table 1 provides an example of base values $B$, grouped by size class, the ratios $\epsilon/B$, which are considered to be outliers, and the ranges of $\epsilon=|F-B|$. The midpoints of the $B$, and $\epsilon$ ranges are also shown, for later use. Table 1 shows some peculiarities. The ranges corresponding to $\epsilon/B=1.50$ and 3.00 have the same minimum $\epsilon$, effectively violating Assumption 3. Worse, the range corresponding to $\epsilon/B=.40$ has all of its $\epsilon$ range higher than the corresponding values in the adjacent ranges. This is in spite of a reasonable progression of values of $\epsilon/B$, which accords with Property 1. One solution to these problems is to have the decision-maker revise the $B$ size classes and/or the values of $\epsilon/B$. However, this may not always be possible. In these cases, the estimation technique for $\ell$ described below must be used with the violating ranges thrown out.
Table 1: $B$ and $\epsilon$, by Size Class
@@ 3.2 Preexisting Outlier Criteria
The two rightmost columns of Table 1 show the midpoints of the ranges of $B$ and $\epsilon$. These are used to construct $\ell$ by linear regression. Essentially, these points are used to define a level curve of $\ell$ in the $(B,\epsilon)$ space. By the Implicit Function Theorem, these points satisfy the equation
$\frac{d\epsilon}{dB}=-\frac{\partial\ell/\partial B}{\partial\ell/\partial\epsilon}=-\frac{q\epsilon B^{q-1}}{B^{q}}=-\frac{q\epsilon}{B}.$ (17)
Upon rearrangement, equation (17) becomes
$\frac{d\epsilon}{\epsilon}=-q\frac{dB}{B}.$ (18)
@@
$\log\epsilon=-q\log B+K.$ (19)
where $K$ is a constant of integration. Equation (19) can then be estimated by linear regression to find the values of the parameters –q and K. (19) is then exponentiated and rearranged to obtain
$\ell(\epsilon,B)=\epsilon B^{q}=e^{K}=C.$ (20)
@@ 3.3 GIS
Population change between censuses is of great interest. The main problem in interpreting change is addressed by Property 1: The coefficient of variation decreases in population. We examine population change by county 2010-2020. The 2010 population is the U.S. Census Bureaus’ Population Estimates Program 2010 base. The 2020 population is the population as enumerated by the 2020 census. The Maps omit Alaska and Hawaii for convenience.
@@
Map 1 displays absolute changes in population. This map has some features of note. The first set consists of groups of low growth The first of these is the band of counties in the Western Plains with low values. The second is a barely discernable band of counties inland of the Fall Line from Pennsylvania to Virginia. Finally, another band of counties with low values runs along the northern borders of Tennessee and Arkansas. All of these are interrupted by high value counties. Their values generally reflect their low populations. The Washington-Boston megalopolis is easily seen with its high absolute differences. This megalopolis really extends from Richmond, Virginia to Portland, Maine.
@@
Map 2 displays the absolute percent differences. The first thing we notice is the lesser clustering of counties relative to Map 1. The Western Plains counties identified earlier show a great range of absolute growth rates, reflecting their small sizes. The band of low growth counties inland of the Fall Line has grown denser and longer, reaching into Tennessee and parts of New England. The high absolute difference areas of Map 1 also generally have high absolute percentage growth rates: More urban counties are experience higher absolute growth rates. The Boston-Washington metropolis splits into 4 areas: an area of high absolute percentage growth from Richmond to Washington, an area of lesser absolute percentage growth around Boston and Portland, an area of even less absolute percentage growth from Baltimore to New York City and an area of very low absolute percentage growth in central Connecticut.
@@
Map 3 displays the q=.5q=-.5 loss functions by county. The low growth area in the Western Plains expands to Wyoming, Minnesota, Michigan, Iowa and Missouri. While many of the counties have darker coloration than in Map 1, that could be an artifact of the ranges. The lines of low growth counties in the East remain, but are thinner. The Boston-Washington megalopolis reappears with the exception of central Connecticut. By varying $q$ and the ranges, the analyst gain insights into county population changes. Using a signed loss function would enable identifying growing and declining counties separately and provide better insights. Doing these enables loss functions to become data exploration tools.
3.4 Example with Outliers Classified by Another Variable
@@
This example contains outlier criteria, $D$, for the variable of interest $V$, classified by another variable, [^10] The authors would like to thank Gregg R. Diffendal for initial guidance with this example. $R$, the reference variable. DDcan represent differences or ratios, or more complicated functions, of successive values of $V$. Normally, the outlier criteria $D$ decline in $R$, so that changes in $V$ take on less importance as $R$ rises. Table 2 provides an example of a set of these criteria.
Table 2: Values of $R$ and $D$
@@
Figure 1 plots $D$ as a function of $R$, where $R$ is taken to be the minimum value of its range and both variables are plotted on logarithmic axes. Points lying above this curve are considered to be outliers.
Figure 1: $D$ as a Function of $R$
Figure 1 shows that the $(R,D)$ curve is almost a straight line in logarithms. Therefore, we are led to statistically estimate the equation
$\log D=a+b\log R.$ (21)
The resulting equation will produce a straight line intersecting the $(R,D)$ curve, when viewed in logarithms. Upon rearrangement, we obtain the equation of criticality
$DR^{-b}=e^{a}=C.$ (22)
An outlier is declared whenever $DR^{-b}>C$. Using the points plotted in Figure 1, the estimated coefficients are $a=4.889506$ and b=0.33692b=-0.33692. Substituting these into equation (22) obtains the estimated equation of criticality
$DR^{0.33692}=132.8879.$ (23)
This is similar to the loss functions described in Section 2 and estimated in Subsection 3.2, with the important difference that the exponent of $R$ is positive. This occurs because $D$ and $R$ are inversely correlated and $D$ is not an explicit function of $R$.
In the authors’ experience, equation (23) resulted in missing too many outliers. [^11] This is an example of the empirical nature of loss functions. We decided, instead, to draw a straight line (in logarithms) between the two endpoints, so as to capture all of the previously declared outliers. The resulting equation (including a slight adjustment to $b$ to ease interpretation) is
$DR^{\frac{1}{3}}=80.$ (24)
@@
4 Conclusion

### v2

@@ I.3.1 General Procedure for Using Loss Functions
I.3.2 Preexisting Outlier Criteria
Table I.1 provides an example of base values $B$, grouped by size class, the ratios $\epsilon/B$, which are considered to be outliers, and the ranges of $\epsilon=|F-B|$. The midpoints of the $B$, and $\epsilon$ ranges are also shown, for later use. Table I.1 shows some peculiarities. The ranges corresponding to $\epsilon/B=1.50$ and 3.00 have the same minimum $\epsilon$, effectively violating Assumption I.3. Worse, the range corresponding to $\epsilon/B=.40$ has all of its $\epsilon$ range higher than the corresponding values in the adjacent ranges. This is in spite of a reasonable progression of values of $\epsilon/B$, which accords with Property 1. One solution to these problems is to have the decision-maker revise the $B$ size classes and/or the values of $\epsilon/B$. However, this may not always be possible. In these cases, the estimation technique for $\ell$ described below must be used with the ranges whose midpoints violate monotonicty discarded.
Table I.1: $B$ and $\epsilon$, by Size Class
@@ I.3.2 Preexisting Outlier Criteria
The two rightmost columns of Table I.1 show the midpoints of the ranges of $B$ and $\epsilon$. These are used to construct $\ell$ by linear regression. Essentially, these points are used to define a level curve of $\ell$ in the $(B,\epsilon)$ space. By the Implicit Function Theorem, these points satisfy the equation
$\frac{d\epsilon}{dB}=-\frac{\partial\ell/\partial B}{\partial\ell/\partial\epsilon}=-\frac{q\epsilon B^{q-1}}{B^{q}}=-\frac{q\epsilon}{B}.$ (I.17)
Upon rearrangement, equation (I.17) becomes
$\frac{d\epsilon}{\epsilon}=-q\frac{dB}{B}.$ (I.18)
@@
$\log\epsilon=-q\log B+K.$ (I.19)
where $K$ is a constant of integration. Equation (I.19) can then be estimated by linear regression to find the values of the parameters –q and K. (I.19) is then exponentiated and rearranged to obtain
$\ell(\epsilon,B)=\epsilon B^{q}=e^{K}=C.$ (I.20)
@@ I.3.3 GIS
Population change between censuses is of great interest. The main problem in interpreting change is addressed by Property I.1: The coefficient of variation decreases in population. We examine population change by county 2010–2020. The 2010 population is the U.S. Census Bureaus’ Population Estimates Program 2010 base. The 2020 population is the population as enumerated by the 2020 census. The Maps omit Alaska and Hawaii for convenience.
@@
Map I.1 displays absolute changes in population. This map has some noteworthy features. The first set consists of groups of low growth The first of these is the band of counties in the Western Plains with low values. The second is a barely discernable band of counties inland of the Fall Line from Pennsylvania to Virginia. Finally, another band of counties with low values runs along the northern borders of Tennessee and Arkansas. All of these are interrupted by high value counties. Their values generally reflect their low populations. The Washington-Boston megalopolis is easily seen with its high absolute differences. This megalopolis really extends from Richmond, Virginia to Portland, Maine.
@@
Map I.2 displays the absolute percent differences. The first thing we notice is less clustering of counties relative to Map I.1. The Western Plains counties identified earlier show a great range of absolute growth rates, reflecting their small sizes. The band of low growth counties inland of the Fall Line has grown denser and longer, reaching into Tennessee and parts of New England. The high absolute difference areas of Map I.1 also generally have high absolute percentage growth rates: More urban counties are experience higher absolute growth rates. The Boston-Washington metropolis splits into 4 areas: an area of high absolute percentage growth from Richmond to Washington, an area of lesser absolute percentage growth around Boston and Portland, an area of even less absolute percentage growth from Baltimore to New York City and an area of very low absolute percentage growth in central Connecticut.
@@
Map I.3 displays the q=.5q=-.5 loss functions by county. The low growth area in the Western Plains expands to Wyoming, Minnesota, Michigan, Iowa and Missouri. While many of the counties have darker coloration than in Map I.1, that could be an artifact of the ranges. The lines of low growth counties in the East remain, but are thinner. The Boston-Washington megalopolis reappears with the exception of central Connecticut. By varying $q$ and the ranges, the analyst can gain insights into county population changes. Using a signed loss function would enable identifying growing and declining counties separately and provide better insights. Doing these enables loss functions to become data exploration tools.
I.3.4 Example with Outliers Classified by Another Variable
@@
This example contains outlier criteria, $D$, for the variable of interest $V$, classified by another variable, [^10] The authors would like to thank Gregg R. Diffendal for initial guidance with this example. $R$, the reference variable. $D$ can represent differences or ratios, or more complicated functions, of successive values of $V$. Normally, the outlier criteria $D$ decline in $R$, so that changes in $V$ take on less importance as $R$ rises. Table I.2 provides an example of a set of these criteria.
Table I.2: Values of $R$ and $D$
@@
Figure I.1 plots $D$ as a function of $R$, where $R$ is taken to be the minimum value of its range and both variables are plotted on logarithmic axes. Points lying above this curve are considered to be outliers.
Figure I.1: $D$ as a Function of $R$
Figure I.1 shows that the $(R,D)$ curve is almost a straight line in logarithms. Therefore, we are led to statistically estimate the equation
$\log D=a+b\log R.$ (I.21)
The resulting equation produces a straight line intersecting the $(R,D)$ curve, when viewed in logarithms. Upon rearrangement, we obtain the equation of criticality
$DR^{-b}=e^{a}=C.$ (I.22)
An outlier is declared whenever $DR^{-b}>C$. Using the points plotted in Figure I.1, the estimated coefficients are $a=4.889506$ and b=0.33692b=-0.33692. Substituting these into equation (I.22) obtains the estimated equation of criticality
$DR^{0.33692}=132.8879.$ (I.23)
This is similar to the loss functions described in Section I.2.1 and estimated in Subsection I.3.2, with the important difference that the exponent of $R$ is positive. This occurs because $D$ and $R$ are inversely correlated and $D$ is not an explicit function of $R$.
In our experience, equation (I.23) resulted in missing too many outliers. [^11] This is an example of the empirical nature of loss functions. We decided, instead, to draw a straight line (in logarithms) between the two endpoints, so as to capture all of the previously declared outliers. The resulting equation (including a slight adjustment to $b$ to ease interpretation) is
$DR^{\frac{1}{3}}=80.$ (I.24)
@@
Part II Data of Mixed Sign
II.1 Introduction to Part II
We remove the restriction on the direction of the signs in this Part. This necessitates the creation of much machinery to create useful loss functions. In Section II.2) we specifiy the assumptions and develop the simplest loss function.develops the signed loss function, which incorporates the direction of the difference. In Section II.2.2 we developed the signed loss function and use it to find the optimal parametrization. Unlike Part I, the parametrization depends on the geometry of the loss function, rather than distributions of the data series.
II.2 The Loss Function
We describe the assumptions and create the simplest loss function, $L(F,B)$, where $F$ is the future value and $B$ is the base period value. Time can be either chronological or nominal, in which $F$ and $B$ are simply different sets of data. The loss function is the penalty associated with the difference between $F$ and $B$. Roughly speaking, the greater the difference between $F$ and $B$, the greater the loss. After the necessary assumptions are made, the simplest form of $L$ is specified. Restrictions on the values of the parameters of $L$ which make it increase in $B$ for a given average absolute relative difference are then specified. We give the assumptions and construct the basic form of the unsigned loss function in Subsection II.2.1. We prove that $(0,0)$ is a continuity point of $L$ under a reasonable assumption in Subsubsection II.2.1.2. Because of the need to deal with the absolute values in $B$ and $F$, we construct the average absolute relative difference and use it to fully specify $L$ in Subsubsection II.2.1.1. We specify the outlier criteria in Subsection II.2.1.3.
II.2.1 The Unsigned Loss Function
The unsigned loss function $L$ is constructed by making five assumptions. The first assumption is that $L$ is defined everywhere in the real plane $\Re^{2}$:
Assumption II.1 (unrestricted domain): For all $(F,B)\in\Re^{2}$, $L(F,B)$ is defined and single-valued.
This and Assumption II.2 generalize Assumption I.1.
Assumption II.2 makes $L$ symmetric in the difference $\epsilon$, whether applied to $B$ or $F$:
Assumption II.2 (symmetry in differences): $L(B+\epsilon,F)=(B-\epsilon,F)$ and $L(B,F+\epsilon)=(B,F-\epsilon)$ for all $B,F,\epsilon>0$.
This assumption is relaxed in Section II.2.2.
A desirable property is that $L$ be symmetric with respect to its arguments. To give a concrete example, we want L(1,1000)=L(1000,1)L(-1,1000)=L(1000,-1). This stated formally as Assumption II.3:
Assumption II.3 (symmetry in arguments): $L(B,F)=L(F,B)$.
We now introduce some new notation. Let $X=|F|$ and $Y=|B|$. Let the loss function $\ell(\epsilon,\Sigma)\equiv L(F,B)$, where $\epsilon=F-B$ and $\Sigma=\Sigma(X,Y)$ is a function such that $\partial\Sigma/\partial X>0$ and $\partial\Sigma/\partial Y>0$ . Assumption II.3 implies that $\Sigma(X,Y)=\Sigma(Y,X)$, so that $\Sigma$ is symmetric in its arguments. The remaining Assumptions are stated in terms of $\ell$.
Assumption II.4 makes $\ell$ (and $L$) increasing in the difference $\epsilon$:
Assumption II.4 (monotonicity in difference): $\partial\ell/\partial\epsilon>0$ for $\epsilon\geq 0$.
This is identical to Assumption I.2.
Finally, we want $\ell$ to decrease in $\Sigma$.
Assumption II.5 (monotonicity in arguments): $\partial\ell/\partial\Sigma<0$ for all $\Sigma\geq 0$.
This is identical to Assumption I.3.
The simplest function which satisfies Assumptions II.1–II.5 is [^12] An infinite number of loss functions satisfy Assumptions 1–5. This one is merely the simplest.
$\ell(\epsilon,\Sigma)=\left\{\begin{array}[]{cc}\epsilon^{p}\Sigma^{q}&\Sigma>0\\ 0&\Sigma=0,\end{array}\right$. (II.1)
where $p>0$ and $q<0$. Note that equation (II.1) is stated in terms of $\epsilon$ and $\Sigma$. The simplest form of $\Sigma$ will be determined in equation (II.2) below. Theorem II.1 below shows that setting $\ell(0,0)=0$ makes $L$ continuous at $(0,0)$, given the choices of $\Sigma,p,q$. This way of defining $\ell(0,0)$ avoids problems associated with division by 0.
II.2.1.1 Determination of $\Sigma$ and $L$
From equation (II.1), it is clear that $\ell(0,\Sigma)>0$ for all $\Sigma>0$. We would like to define $\Sigma$ so that whenever eitherX>0X>0 or $Y>0,\Sigma>0.$ We would also like $\Sigma(0,0)=0$. The simplest equation for $\Sigma$ is
$\Sigma=X+Y=|F|+|B|.$ (II.2)
Substituting equation (II.2) into equation (II.1) and expanding $\Sigma$ obtains the equation for $L$:
$L(F,B)=\left\{\begin{array}[]{cc}|F-B|^{p}\big(|F|+|B|\big)^{q}&B\text{ or }F\neq 0\\ 0&B=F=0.\end{array}\right$. (II.3)
Setting $L(0,0)=0$ is a convention to simplify its definition. Subsubsection II.2.1.2 shows that this is consistent with $(0,0)$ being a limit point for $L$ when $p+q>0$. s
II.2.1.2 Proof that $(0,0)$ is a Continuity Point of $L$ when $p+q>0$
Theorem II.1: $L(0,0)=0$ is a continuity point of $L(F,B)$ when p+q>0p+q>0.131313This theorem was proven with the help of ChatGPT.
Proof: Let $Z=|F|+|B|$ and $r=|F-B|/(|F|+|B|)\in[0,1]$, taking r = 0 at the origin. Then,
$L=Z^{p+q}r^{p}.$ (II.4)
Hence, $0\leq L\leq Z^{p+q}$. Since $p+q>0$ and $Z\rightarrow 0$ as $F,B\rightarrow 0$, $L\rightarrow 0$ as $F,B\rightarrow 0$ by the squeeze theorem. Therefore, $(0,0)$ is a continuity point of $L.\Box$
Note that the proof fails for $p+q\leq 0$, where no finite limit exists.
II.2.1.3 Outlier Criteria
See Section I.2.1 for a discussion of outlier criteria and the Lie symmetry that enables the use of $p=1$. Thus,
$L(F,B)=\left\{\begin{array}[]{cc}|F-B|\big(|F|+|B|\big)^{q}&B\text{ or }F\neq 0\\ 0&B=F=0.\end{array}\right$. (II.5)
A desirable property of the loss function is that it rises in |FB||F-B| for a given average absolute relative difference. The problem lies in defining this difference. The absolute relative differences with respect to $|F|$ and $|B|$ are $|F-B||F|^{-1}$ and $|F-B||B|^{-1}$, respectively. Their harmonic mean is
$2|F-B|(|F|+|B|)^{-1}.$ (II.6)
Removing the 2 as a nuisance constant, we define the average absolute percentage difference as
$|F-B|\big(|F|+|B|\big)^{-1}=\epsilon\Sigma^{-1}$ (II.7)
for $\Sigma>0$. Equation (II.7) is the first row of equation (II.5) with q=1q=-1. Choosing q>1q>-1 makes the loss function rise in $|F|+|B|$, for a given average absolute relative difference. This is also required by Theorem II.1. We state this as Property II.1:
Property II.1: The loss function defined by equations (II.5) increases in $|F|+|B|$ for any given average absolute percentage difference. This is assured whenever q>1q>-1.
The reader may note that $q=0$ turns the first row of equation (II.5) into the absolute values of the differences. Thus, values of $q$ between 0 and 1-1 represent various tradeoffs of absolute differences and average absolute percentage differences. Consider the product of the $r$th power of the absolute difference and the $s$th power of the average absolute relative difference, where $r,s>0$
$|F-B|^{r}\left(\frac{|F-B|}{|F|+|B|}\right)^{s}$ (II.8)
This function is isomorphic to the function $|F-B|\big(|F|+|B|\big)^{r/(r+s)}$. Thus, these intermediate values of $q$ correspond to an infinite number of pairs $(r,s)$ where $q=r/(r+s)$. Geometrically, the same loss function is generated for all pairs $(r,s)$ lying on the line s=(11/q)rs=-(1-1/q)r.
II.2.2 The Signed Loss Function
At times, not only the value of the loss function is important, but also the sign of the difference. Different outlier generation processes may manifest themselves by producing predominantly positive or negative differences. We can account for these by creating the signed loss function $S$, which is simply the loss function $L$, multiplied by the signum function of the difference
$S(F,B)=\left\{\begin{array}[]{cc}|F-B|\big(|F|+|B|\big)^{q}\operatorname{sgn}(F-B)=(F-B)\big(|F|+|B|\big)^{q}&B\text{ or }F\neq 0\\ 0&B=F=0,\end{array}\right$. (II.9)
where $\operatorname{sgn}x=+1$ for $x>0$, 0 for $x=0$, and 1-1 for $x<0$.
Using $S$, one can create different critical values for loss, depending on whether the difference is positive or negative. To wit, one can pick $C_{+},C_{-},C+\neq-C_{-}$, such that a pair $(F,B)$ is declared an outlier if either $S(F,B)<C_{-}$ or $S(F,B)>C_{+}$. Again, the choice of whether to use $S$ and then use asymmetric critical bounds is an empirical matter. [^14] The asymmetry need not be limited to the critical values. The signed loss function can incorporate different values of $q$, depending on the sign of the difference. The next Subsection investigates $S$ graphically to study the behavior of these loss functions and to suggest choices for $q$. We can see that the geometry is very important to this choice.
II.2.3 Behavior of $S$ and the Choice of $q$
The behavior of $S$ requires that plots of it for various values of $q$ be examined to obtain a reasonable loss function. The limiting functions when $q=0$ and q=1q=-1 are of interest. $q=0$ implies that S(F,B)=FBS(F,B)=F–B. This defines a plane in $\Re^{3}$, as shown in Figure II.1.
[Insert Figure II.1 about here]
Figure II.2 illustrates $S$ when q=1q=-1.
[Insert Figure II.2 about here]
Figure II.2 shows the lack of a limit at (0,0). Later Figures have similar artifactual white lines. Another thing is very apparent from this graph, whenever $B$ and $F$ are of opposite signs, $S(F,B)=\pm\operatorname{sgn}F$. This can be seen by substituting q=1q=-1 into equation (II.9) when $B$ or $F$ is nonzero:
$S(F,B)=\frac{F-B}{|F|+|B|}$ (II.10)
Noting that $|x|=x$ when $x>0$ and |x|=x|x|=-x when $x<0$, we can examine the behavior of $S$ when $B$ and $F$ are of opposite signs. When $F>0$ and $B<0$, equation (II.10) becomes
$S(F,B)=\frac{F-B}{|F|+|B|}=\frac{F-B}{F-B}=1=\operatorname{sgn}F$ (II.11)
The reader may verify that $S(F,B)=-1=\operatorname{sgn}F$ when $F<0$ and $B>0$. When $F>0$ and $B>0$, the limit becomes path-dependent, lying on $[0,1]$. Finally, when $F<0$ and $B<0$, the limit lies on [1,0][-1,0], with a similar path-dependence. The proofs are omitted. Thus, instead of $S$ having a limit at $(0,0)$, it has a limit set of [1,1][-1,1].
Figure II.2 shows cusps along the axes. These are also apparent in Figures II.3II.11, which graph $S$ for $q$ ranging from .1-.1 to .9-.9 in increments of .1-.1, respectively.
[Insert Figures II.3II.11 about here]
Figures II.3II.5 are almost planar, since $q$ is close to 0. However, Figure II.5, q=.3q=-.3, begins to show the flattening out behavior of Figure II.2. This flattening out increases as $q$ declines to .9-.9 in Figure II.11. In this Figure, $S\approx\operatorname{sgn}F$ when $B$ and $F$ are of opposite signs.
Given the nonsmooth behavior shown in Figures II.3II.11, the analyst has to decide on a value of $q$ which produces reasonable behavior. It appears that intermediate choices of $q$ are best behaved: these offer a good compromise between simply taking the difference between $F$ and $B$ ($q=0$) and generating the flatness in $S$ when $q$ approaches 1-1. Similar to Part I, we recommend starting with q=.5q=-.5.
II.3 Figures
Figure II.1: $S(F,B),q=0$
Figure II.2: S(F,B),q=1S(F,B),q=-1
Figure II.3: S(F,B),q=.1S(F,B),q=-.1
Figure II.4: S(F,B),q=.2S(F,B),q=-.2
Figure II.5: S(F,B),q=.3S(F,B),q=-.3
Figure II.6: S(F,B),q=.4S(F,B),q=-.4
Figure II.7: S(F,B),q=.5S(F,B),q=-.5
Figure II.8: S(F,B),q=.6S(F,B),q=-.6
Figure II.9: S(F,B),q=.7S(F,B),q=-.7
Figure II.10: S(F,B),q=.8S(F,B),q=-.8
Figure II.11: S(F,B),q=.9S(F,B),q=-.9
Part III Conclusion